OpenStudy (anonymous):

How would I explain that my equation is has a solution of (-1, 5). And I already know that it is (-1,5). y=2x+7 y=x+6

Nnesha (nnesha):

replace x and y in the equation by your solution if both side equal then that is ur right answer

OpenStudy (lukaryse):

First in, plug it the x value. We know that (-1,5) has the same form of (x,y). So we just need to plug in the x value to evaluate it. y=2x+7 y=2(-1)+7 y=-2+7 Y=5 y=x=6 y=-1+6 Y-5

OpenStudy (anonymous):

you have to substitute one equation into the other hint, y is easier to sub than x is

Nnesha (nnesha):

 solve right side if you get 5 on right side then yes your answer is right